OpenStudy (shealah):

XX thinks that is you increase a number by 10% and then decrease it by 10% you are right back where you started. Prove XX is incorrect by using a specific example?

OpenStudy (anonymous):

Let the number be 100 10% of 100 = (10/100)* 100 = 10 So new increased number = 100 + 10 = 110 Now 10% of 110 = (10/100)*110 = 1100/100 = 11 So on subtracting we get 110 - 11 = 99 So we end up with a number which is less than the 100 we started with..